A 9.0-V battery is connected to a bulb whose resistance is 3.0 Ω .

Part A

How many electrons leave the battery per minute?

Express your answer using two significant figures.
Answer #1

here,

Voltage, v = 9 V
Resistance, R = 3 Ohms

From ohms law, Current = Volatge/Resistance
I = V/R
I = 9/3
I = 3 A

per minute
I = 3 *60 = 180 C leaving the wire

Therefore
No of electron per minute will = 180/(1.6*10^-19) = 1.125*10^21